OpenStudy (anonymous):

Given f(x) = the quantity of 3x minus 1, divided by 2 solve for f-1(4). Please help

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@surd

OpenStudy (anonymous):

@gudden

OpenStudy (anonymous):

?

OpenStudy (anonymous):

>>>>>>>????????????

OpenStudy (anonymous):

its an ambush

OpenStudy (gudden):

ahahah xd..:p @LastDayWork . u also come n.. join in here :p

OpenStudy (anonymous):

its an ambush irepeat its an ambush

OpenStudy (anonymous):

bob is hit

OpenStudy (anonymous):

soldiers secure the perimeter

OpenStudy (anonymous):

wtf their are enemy on the front line

OpenStudy (anonymous):

firing from west take cover

OpenStudy (anonymous):

lol kngiht is here

OpenStudy (anonymous):

help is on its way

OpenStudy (lastdaywork):

OKAY FOLKS.....PARTY'S OVER DELETE ALL THESE SPAMS....OR GET REPORTED

OpenStudy (gudden):

@LastDayWork ..so u r newspaper reporter,.. :p

OpenStudy (lastdaywork):

Please delete all unnecessary posts.. Take this RPG on OS Feedback..

OpenStudy (gudden):

haha u killed me... my dear V :p

OpenStudy (lastdaywork):

@Imbueds Please clarify what do you mean by " f-1(4) "

OpenStudy (lastdaywork):

Try not to hijack a new-comer's feed.. :P

OpenStudy (anonymous):

\[f \left( x \right)=\frac{ 3x-1 }{2 }\] let f(x)=y \[y=\frac{ 3x-1 }{ 2 }\] interchange x and y \[x=\frac{ 3y-1 }{2 },2x+1=3y,y=\frac{ 2x+1 }{3 },f ^{-1}\left( x \right)=\frac{ 2x+1 }{ 3 }\] put x=4 and get the solution.

OpenStudy (anonymous):

thank you Surji

OpenStudy (anonymous):

my name is surjit (pronounced as surjeet)

OpenStudy (anonymous):

yw